Keith weighs 20kg more than Beth, while Henry weighs 30kg less than twice as much as Beth. If Keith and Henry weigh the same,how much does Beth weigh(in kg)?
Found 2 solutions by mukhopadhyay, priya:
Answer by mukhopadhyay(490)

Let Beth's weight be x kg
Keith's weight = x+20 kg
Henry's weight = 2x-30 kg;
Keith's weight = Henry's weight
=> x+20 = 2x-30
=> x = 50;
Answer: Beth's weight is 50 kg.

Answer by priya(51)

First try to write the given informations in the form of equations.
1. let K be keith's weight ,H be henry's weight , B be beth's weight.

then u can write the given infy's as,
eq 1. K = 20 + B ,
eq 2. H = 2 * B -30,
K & H ARE SAME K = H,
Therefore eq 1. and eq 2. are same so they can be equated,
20 + B = 2* B - 30
20 + 30 = 2* B - B ( RE-ARRANGE NUMBERS ON ONE SIDE AND VARIABLE "B" ON OTHER SIDE ,u've to change the sign when doing so)
Therefore B = 50, The weight of BETH is 50kg..
WANT TO LEARN ONLINE..VISIT MY SITE..
PRIYA